# SRLM: Human-in-Loop Interactive Social Robot Navigation with Large Language Model and Deep Reinforcement Learning

### v2

@@ V-A5 Effectiveness of LNM
Figure 6: The box plot of social score SS.
From Fig. 6, the ablation model SR-RLNM achieves a $90\%$ SR and $90$ SS, which is slightly lower than the performance of SRLM overall. We can observe that SR-RLNM can provide a benchmark performance derived from the RLNM capability. However, as shown in Fig. 7, SR-RLNM demonstrates a similar path quality to SRLM in the early timesteps, but SR-RLNM still adheres to previous strategic preferences even after receiving renovation feedback from the user updating the task objective and personal preference. Hence, the introduction of LNM in SRLM stabilizes the robustness of navigation performance, especially for the stage after user feedback, because the pre-trained DRL-based policy cannot adjust itself to highly adaptive behavioural representations in the execution stage.
(a) Policy: SR-RLNM
(b) Policy: SRLM
(c) Policy: SR-LNM
(d) Policy: SRLM
(e) Policy: SR-LFM
(f) Policy: SRLM
Figure 7: Comparison of trajectory visualization: Visualization of trajectories for ablation models and SRLM, all tested using the same test case.
@@ V-A6 Effectiveness of RLNM
Despite the amazing inference ability exhibited by large language models in many applications, LLM-based developments of sequential control systems have struggled with highly dynamic environments and sequential data dimensions. The environments of social navigation tasks require robots to understand potential pedestrian intents and engage in HRI cooperation to adapt to environmental dynamics. Such challenges, coupled with the lack of RLNM compared to SRLM, result in struggling and precarious effects in dynamic scenarios. From both the average results and trajectory instances shown Fig. 7, we observe that SR-LNM exhibits many instances of reciprocal dance phenomena, where the robot swings from side to side. Thus, to maintain benchmarking performance for LLM-based social robots in dynamic spaces, we recommend continued use of DRL-based robot executors.
@@ V-A7 Effectiveness of LFM
As observed in Fig. 6 and Fig. 7, we find that SR-LFM exhibits limited planner capability compared to SRLM. The blunt fusion with fixed relative weights presents lower flexibility than LFM, because the LLM-driven LFM can infer the situation for a better fusion strategy from more evidence chains with the developments of GoT. The blending mechanism of LFM is significant for adjusting and fusing two robot actions, leveraging the inference ability from LLM.
@@ VI Conclusion
In this work, we developed an interactive social robot large model. SRLM leverages the inference ability of LLM to interpret user language commands and enhance the adaptability of DRL-based navigation policy. Additionally, the GoT is developed by LFM to evaluate the relative action score of the executed actions from LLM-based and DRL-based planners, incorporating LNM and RLNM blocks. Finally, SRLM demonstrates outstanding efficiency compared to baselines and ablation models in both simulation and real-world experiments.

### v3

@@ V-A5 Effectiveness of LNM
From Table I, the ablation model SA-RLNM achieves a $75\%$ SR and $74$ SS, which is slightly lower than the performance of SALM overall. We can observe that SA-RLNM can provide a benchmark performance derived from the RLNM capability. However, as shown in the common trajectory maps in the experiment, SA-RLNM demonstrates a similar path quality to SALM in the early timesteps, but SA-RLNM still adheres to previous strategic preferences even after receiving renovation feedback from the user updating the task objective and personal preference. Hence, the introduction of LNM in SALM stabilizes the robustness of navigation performance, especially for the stage after user feedback, because the pre-trained DRL-based policy cannot adjust itself to highly adaptive behavioral representations in the execution stage.
@@ V-A6 Effectiveness of RLNM
Despite the amazing inference ability exhibited by large language models in many applications, LLM-based developments of sequential control systems have struggled with highly dynamic environments and sequential data dimensions. The environments of social navigation tasks require robots to understand potential pedestrian intents and engage in HRI cooperation to adapt to environmental dynamics. Such challenges, coupled with the lack of RLNM compared to SALM, result in struggling and precarious effects in dynamic scenarios. From both the average results and common trajectory maps in the experiment, we observe that SR-LNM exhibits many instances of reciprocal dance phenomena, where the robot swings from side to side. Thus, to maintain benchmarking performance for LLM-based social robots in dynamic spaces, we recommend continued use of DRL-based robot executors.
@@ V-A7 Effectiveness of LFM
As observed in Table I and common trajectory maps in the experiment, we find that SR-LFM exhibits limited planner capability compared to SALM. The blunt fusion with fixed relative weights presents lower flexibility than LFM, because the LLM-driven LFM can infer the situation for a better fusion strategy from more evidence chains with the developments of GoT. The blending mechanism of LFM is significant for adjusting and fusing two robot actions, leveraging the inference ability from LLM, which can effectively improve overall performance. More experiment videos can be found at website: https://sites.google.com/view/navi-salm.
@@ VI Conclusion
In this work, we developed an interactive social robot large model. SALM leverages the inference ability of LLM to interpret user language commands and enhance the adaptability of DRL-based navigation policy. Additionally, the GoT is developed by LFM to evaluate the relative action score of the executed actions from LLM-based and DRL-based planners, incorporating LNM and RLNM blocks. Finally, SALM demonstrates outstanding efficiency compared to baselines and ablation models in the experiments.